# StyleReiser: Stylizing Video With Reinforced Structure Guide

### v1

@@ 5.5 Direct optimization with the image conditioning prior
$\mathcal{L}_{\text{cSDS}}$
@@
Figure 16: Comparison of a stylization results obtained by the optimization with the lineart loss term $\mathcal{L}_{\text{lineart}}$ (5) and with the lineart-guided $\mathcal{L}_{\text{cSDS}}$ term (3). Increasing the strength of the loss term $\mathcal{L}_{\text{lineart}}$ does not enforce the structural elements from the target frame to the stylized output as in the case of the $\mathcal{L}_{\text{cSDS}}$ term.
In our experiments, we show that our proposed loss term $\mathcal{L}_{\text{cSDS}}$ (3) enforces structures from the target frame $\mathbf{x}$ to the stylized output of our method. The purpose of this experiment is to show that the $\mathcal{L}_{\text{cSDS}}$ loss built on the ControlNet model cannot be replaced by a simpler loss that uses the lineart image conditioning model $c(\mathbf{x})$, which is already part of ControlNet, where it provides the guidance image $\mathbf{c}$. In Fig. 16 we show that replacing the lineart-guided $\mathcal{L}_{\text{cSDS}}$ with the lineart term
$\mathcal{L}_{\text{lineart}}(\mathcal{Z};\mathbf{\Phi},\mathbf{\Theta}_{c})=\frac{1}{N}\sum_{i\not\in\mathbf{K}}\left\|c\left(\hat{\mathbf{y}}_{i};\mathbf{\Theta}_{c}\right)-c\left(\mathbf{x}_{i};\mathbf{\Theta}_{c}\right)\right\|_{2}^{2},$ (5)
where $\mathbf{\Theta}_{c}$ are the parameters of the model $c(\mathbf{x};\mathbf{\Theta}_{c})$ producing the lineart control image $\mathbf{c}$, does not enforce the structures from the target frame $\mathbf{x}_{i}$ to the stylized output $\hat{\mathbf{y}}_{i}$. In this experiment, we trained on a single NVIDIA A100 GPU for $4$ hours, and we selected the model with the lowest total loss.
@@ 6 Limitations
Although our approach represents an improvement over the current state-of-the-art in diffusion-based as well as keyframe-based video stylization methods, we have identified certain limitations in its application.
In style transfer methods, there exists a delicate balance between maintaining fidelity to the features of the style exemplar images and preserving the structural characteristics present in the content that is being stylized. Current state-of-the-art in keyframe-based video stylization mainly emphasizes fidelity to the original features in the style image. In our approach we aim to produce the stylized content aligned both with the style exemplar as well as with the structural characteristics of the target video sequence. Although we enable the user to find a right balance between style and structure using parameters $\lambda_{c}$ and $t$, we acknowledge that in some situations decreasing the fidelity to the style features may be perceived as a potential limitation.
Our method achieves comparable training time as the approach of Futshik et al. Futschik et al. (2021) which is relatively long and thus can limit the possibility to perform interactive updates that are available, e.g., in the method of Texler et al. Texler et al. (2020b). To address this drawback, in future work we plan to combine Texler et al.’s rapid patch-based training strategy with the calculation of VGG loss in a full-frame setting.
In our proposed workflow, a user is expected to select a keyframe that will be used for training. Although certain guidelines can be followed, a mechanism that would enable automatic selection of suitable keyframes could simplify the preparation phase.
@@ 7 Conclusion
We presented a novel keyframe-based video stylization method that maintains visual coherence across distant frames by adapting to changes in scene structure and adhering to important characteristics of a prescribed visual style. This capability not only enhances the quality of stylized sequences, it also eliminates the need for additional correction keyframes, which can be tedious to prepare or inconsistent when created using diffusion-based methods. Our approach serves as a viable alternative to text-based techniques because it mitigates their structural instability and allows users to make additional modifications to the generated keyframes. Moreover, its real-time inference capability enables interactive applications such as consistently stylized video calls, which are difficult to achieve using diffusion-based approaches.

### v2

@@ 5.5 Direct optimization with the image conditioning prior
$\mathcal{L}_{\text{structure}}$
@@
Figure 17: Comparison of stylization results using the line art loss term $\mathcal{L}_{\text{lineart}}$ (5) versus our line art-guided structure loss term $\mathcal{L}_{\text{structure}}$ (4). Increasing the strength of $\mathcal{L}_{\text{lineart}}$ fails to effectively transfer structural elements from the target frame to the stylized output, whereas $\mathcal{L}_{\text{structure}}$ successfully preserves the target’s structural details.
In our experiments, we show that our proposed loss term $\mathcal{L}_{\text{structure}}$ (4) enforces structures from the target frame $\mathbf{y}$ to the stylized output $\hat{\mathbf{y}}$ of our method. The purpose of this experiment is two fold: first, to demonstrate that the $\mathcal{L}_{\text{structure}}$ loss based on the ControlNet model cannot be replaced by a simpler loss that uses the line art image conditioning model $c(\mathbf{y})$, which is already integrated into ControlNet as the conditioning $c$; and second, to present this simpler loss as an alternative structure-preserving mechanism, serving as a baseline for comparison. In Fig. 17, we show that replacing our proposed line art-guided structure loss $\mathcal{L}_{\text{structure}}$ with the line art term
$\mathcal{L}_{\text{lineart}}(\mathcal{Y})=\frac{1}{|\mathcal{Y}|}\sum_{i\not\in\mathbf{K}}\left\|c\left(\hat{\mathbf{y}}_{i}\right)-c\left(\mathbf{y}_{i}\right)\right\|_{2}^{2}$ (5)
fails to enforce the transfer of structural details from the target frames $\mathbf{y}_{i}$ to the stylized outputs $f(\mathbf{y}_{i})=\hat{\mathbf{y}}_{i}$. In this experiment, we trained on a single NVIDIA A100 GPU for $4$ hours and selected the model with the lowest total loss.
@@ 6 Limitations
Although our approach represents an improvement over the current state-of-the-art in both diffusion-based and keyframe-based video stylization methods, we have identified certain limitations in its application.
In style transfer methods, there exists a delicate balance between maintaining fidelity to the features of the style exemplar images and preserving the structural characteristics present in the content that is being stylized. The current state-of-the-art in keyframe-based video stylization mainly emphasizes fidelity to the original features in the style image. Our approach aims to produce stylized content that aligns both with both the style exemplar and the structural characteristics of the target video sequence. Although we enable users to find the right balance between style and structure using parameters $\lambda_{s}$ and $t$, we acknowledge that in some situations, decreasing the fidelity to the style features may be perceived as a potential limitation.
Our method achieves training times comparable to the approach of Futshik et al. Futschik et al. (2021) which is relatively long and can limit the possibility for interactive updates available in methods like that of Texler et al. Texler et al. (2020b). To address this drawback, in future work we plan to combine Texler et al.’s rapid patch-based training strategy with the calculation of the style loss in a full-frame setting.
In our proposed workflow, the user is expected to select a keyframe that will be used for training. Although certain guidelines can be followed, a mechanism that enables automatic selection of suitable keyframes could simplify the preparation phase.
@@ 7 Conclusion
In this work, we introduced a novel keyframe-based video stylization method that balances the preservation of essential structural elements with adherence to a prescribed visual style. By integrating the line art-guided structure loss term $\mathcal{L}_{\text{structure}}$, our approach overcomes limitations of existing diffusion-based and keyframe-based stylization techniques by preserving structural detail from the input video, enhancing the quality and coherence of stylized sequences, and reducing the need for additional correction keyframes. Real-time inference of our method enables interactive applications such as consistently stylized video calls, which are challenging with existing approaches. By blending style fidelity with structural preservation, our method significantly advances video stylization in both quality and usability.
Appendix A Additional Results and Experiments
In this supplementary material, we first present additional results and comparisons with the current state-of-the-art in diffusion-based video stylization (see Sec. A.1). Four new sequences will be provided (see Figures A.18, A.19, A.20, and A.21) and for two of them we also prepared a comparison with keyframe-based methods (see Figures A.22 and A.23). In addition, we compare our method with the approach of Kim et al. Kim et al. (2024) (see Fig. A.24). Finally, we present results of the real-time video call scenario originally proposed by Texler et al. Texler et al. (2020b) (see Sec. A.2 and Fig. A.25).
A.1 Additional results and comparisons
In this section, we present further comparisons of our method with state-of-the-art video stylization approaches. The results of four additional sequences compared with diffusion-based methods are presented in Figures A.18, A.19, A.20, and A.21. Two sequences are compared with keyframe-based methods in Figures A.22 and A.23. In both cases, additional results support the discussion provided in the main paper (Sec. 4), i.e., it is visible that our approach maintains structural details better than diffusion-based as well as keyframe-based techniques and provides overall stability that was previously difficult to achieve using diffusion-based methods (see our supplementary video).
$\mathbf{y}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure A.18: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization: The target video sequence (see a representative target frame $\mathbf{y}$) has been stylized using diffusion-based approaches (top row): (a) Ceylan et al. Ceylan et al. (2023), (b) Yang et al. Yang et al. (2023), (c) Chu et al. Chu et al. (2024), and (d) Geyer et al. Geyer et al. (2024). One frame from those stylized sequences was used as a keyframe (see small insets). The style of this keyframe has been propagated to the rest of the target sequence $\textbf{y}\in\mathcal{Y}$ using our approach (bottom row). Note how our approach better preserves structural details seen in the target frame. Also, see our supplementary video to compare consistency across the entire sequence. Diffusion-based approaches tend to suffer from notable structural flicker, whereas our approach keeps the structure consistent, yielding considerably more stable results.
$\mathbf{y}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure A.19: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization (cont.): See Fig. A.18 for detailed explanation.
$\mathbf{y}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure A.20: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization (cont.): See Fig. A.18 for detailed explanation.
$\mathbf{y}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure A.21: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization (cont.): See Fig. A.18 for detailed explanation.
target frame $\mathbf{y}$
4 keyframes
1 keyframe
(a) Yang et al. Yang et al. (2023)
(b) Jamriška et al. Jamriška et al. (2019)
(c) Texler et al. Texler et al. (2020b)
(d) Futschik et al. Futschik et al. (2021)
(e) our approach
Figure A.22: Results of our approach in comparison with the state-of-the-art in keyframe-based video stylization: The diffusion-based method of Yang et al. Yang et al. (2023) has been used to generate a stylized sequence (a) from which $4$ keyframes were selected to perform video stylization using methods of Jamriška et al. Jamriška et al. (2019) (b) and Texler et al. Texler et al. (2020b) (c), and $1$ keyframe was selected for the method of Futschik et al. Futschik et al. (2021) (d) and our approach (e). Note how our approach better preserves the structural details seen in the target frame. In our supplementary video, it is also visible that our approach keeps the structure consistent, yielding notably more stable results.
target frame $\mathbf{x}$
2 keyframes
1 keyframe
(a) Geyer et al. Geyer et al. (2024)
(b) Jamriška et al. Jamriška et al. (2019)
(c) Texler et al. Texler et al. (2020b)
(d) Futschik et al. Futschik et al. (2021)
(e) our approach
Figure A.23: Results of our approach in comparison with the state-of-the-art in keyframe-based video stylization: Diffusion-based method of Geyer et al. Geyer et al. (2024) has been used to generate a stylized sequence (a). See Fig. A.22 for detailed explanation.
In Fig. A.24, we present a comparison between our approach and the method of Kim et al. Kim et al. (2024). In the case of the sequence shown in the left half of the figure, our approach is better aligned with the structure of the target frame $\mathbf{y}$. For the sequence in the right half of the figure, the result is not as distinctive, therefore we ask the reader to see our supplementary video where it is visible that our method clearly produces more temporally stable results.
We would like disambiguate the distinction between the conditioning used in our method and the one used by Kim et al. Kim et al. (2024). In our method, we use the prior of a pre-trained text-to-image network conditioned by the guidance image $\mathbf{c}$ to regularize the loss term $\mathcal{L}_{\text{style}}$ (2). By regularizing this term, we enforce the structures from the target frame $\mathbf{y}$ into the stylized frame $\hat{\mathbf{y}}$. We sample the prior of the pre-trained network in every training step and just at a single step $t$ of its denoising schedule. We do not use text conditioning. In contrast, the work of Kim et al. Kim et al. (2024) is built on the method of Brooks et al. brooks_instructpix2pix_2023 that uses two-fold conditioning: the input image and the text instruction. Each conditioning has a guidance scale with which the degree of similarity between the generated samples and the input image is balanced, as well as the degree of similarity with the editing instruction. Both conditions are used during the entire image generation process. The method of Kim et al. Kim et al. (2024) has a significant memory footprint. Therefore, only $90$ stylized frames are available for each sequence featured in our supplementary video.
target frame $\mathbf{x}$
Kim et al. Kim et al. (2024)
our method
target frame $\mathbf{x}$
Kim et al. Kim et al. (2024)
our method
Figure A.24: Comparison with the state-of-the-art in keyframe-based video stylization: Diffusion-based method of Kim et al. Kim et al. (2024) was used to generate a stylized sequence with the edit prompt: “hyperrealistic detailed oil painting of a girl/woman.” From that sequence, $1$ keyframe was selected (see small insets) to perform video stylization using our method. Note how our approach better preserves structural details seen in the target frame. In our supplementary video, you can see how our method outperforms the approach of Kim et al. with respect to the overall structural stability.
A.2 Real-time video call scenario
Texler et al. Texler et al. (2020b) proposed a real-time video call scenario in which the appearance of the call participant is stylized in real time. In this scenario, only a single artist-made keyframe is used to train our method and the methods of Texler et al. Texler et al. (2020b) and Futschik et al. Futschik et al. (2021). In Fig. A.25, we observe a key advantage of our method: By increasing the strength of the loss term $\mathcal{L}_{\text{structure}}$, the structure of the target frame $\mathbf{y}$ is prioritized, and by decreasing its strength, the style characteristics of the style exemplar are emphasized. Compared to the results of Texler et al. Texler et al. (2020b) (a) and Futschik et al. Futschik et al. (2021) (b), our proposed method produces a stylization that adheres better to the structural elements present in the target frame $\mathbf{y}$. To emphasize this fact, we provide three different settings: (c) emphasize more style features of the stylized keyframe, (d) achieve a balance between style and structure, and (e) prioritize structural details present in the target frame.
keyframe
target frame $\mathbf{x}$
Texler et al. Texler et al. (2020b)
Futschik et al. Futschik et al. (2021)
our approach – more style / balanced / more structure
Figure A.25: Our approach applied in the real-time video call scenario originally proposed by Texler et al. Texler et al. (2020b) (a) and later improved by Futschik et al. Futschik et al. (2021) (b) in comparison with three different settings of our approach: (c) emphasizing style features from the stylized keyframe, (d) achieving a balance between style and structure, and (e) prioritizing structure from the target frame. All methods were trained with a single keyframe (left). Note how our approach better preserves the structural details seen in the target frame for all three settings and also has better overall structural stability over the entire sequences as is visible in our supplementary video. The training time of our approach is comparable to the method of Futschik et al. Futschik et al. (2021).